3x^2+8x=9+2x How do I factor and solve using the zero principle?

Answer by ewatrrr(24785)


Hi,
factor and solve using the zero principle
3x^2+8x=9+2x
3x^2+ 6x - 9 =0
factoring
(3x + 9)(x - 1)= 0 Note sum of inner product(9x) and outer product(-3x) = 6x
using the zero principle
(3x + 9)=0
x = -3
(x + 1)= 0
x = -1